# TESSELLATE: Piecing Together the Variable Sky With TESS

### v1

@@ 1 Introduction
The Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2014) has imaged nearly the entire sky in the near-infrared since starting operations in late 2018. Its wide field of view (FoV) of $2160\rm\,deg^{2}$ and its continuous, high-cadence observation strategy have allowed TESS to fill a niche role in the transient community. While initially designed to hunt for exoplanets in the Milky Way, its science use has since expanded widely, including the study of variable stars (Bruch, 2022; Liu & Qian, 2023; Fetherolf et al., 2023), supernovae (SNe, Fausnaugh et al., 2021; Vallely et al., 2021; Fausnaugh et al., 2023; Wang et al., 2023, 2024), gamma-ray burst (GRB) afterglows (Smith et al., 2021; Fausnaugh et al., 2023; Roxburgh et al., 2023; Jayaraman et al., 2024), asteroids (Pál et al., 2018, 2020; Woods et al., 2021; McNeill et al., 2023), comets (Ridden-Harper et al., 2021a; Farnham et al., 2021), exocomets (Zieba et al., 2019; Dobrycheva et al., 2023), stellar flares (Günther et al., 2020; Pietras et al., 2022), and other transient phenomena.
As the TESS mission has evolved, the volume of data produced from each target field, known as a TESS ‘sector’, has greatly increased. Upon launch, it captured full-frame images (FFIs) at a cadence of 30 minutes; this interval was reduced to 10 minutes between 5th July 2020 – 31st August 2022, and has since been further shortened to just 200 seconds.
The data from each sector is released to the public through the MAST archive [^1] http://dx.doi.org/10.17909/0cp4-2j79 (catalog doi:10.17909/0cp4-2j79) as a time-series of FFIs that are ‘calibrated’ but not background subtracted. As such, observers are required to perform their own photometric extraction for sources that are not automatically extracted through the SPOC pipeline (Caldwell et al., 2020). There are a number of successful reduction pipelines available to the public (e.g. Ridden-Harper et al., 2021b; Fausnaugh et al., 2021; Feinstein et al., 2020), each with advantages and disadvantages due to their choice of techniques. However, these pipelines are often small-scale, focusing on producing cutouts of TESS data around desired coordinates or sources. This means that captured information has been left unexamined in the gaps between targeted searches.
@@
In this paper, we introduce TESSELLATE: the TESS Extensive Lightcurve Logging and Analysis of Transient Events pipeline, the first system capable of automatically searching every TESS frame to document all fast transient events captured by TESS. For this pipeline, we use TESSreduce to remove the scattered light background and create difference images. TESSELLATE is purpose-built for use on the OzSTAR supercomputing node held at Swinburne University, and is wrapped in a user-friendly python package [^2] https://github.com/rhoxu/TESSELLATE that allows for simple but thorough control of the various processing steps.
@@ 2.2 Cutting Datacubes
In this stage, each of the generated datacubes are further divided into a number of cutouts, also completed using the astrocut package. This division enables further parallel processing and decreases the size of individual reduction tasks. Furthermore, the TESSreduce package produces more reliable subtractions of the scattered light background for smaller regions with less dynamic range in the background. The number of cutouts is adjustable; by default each datacube is divided into 16 equal cutouts of $512\times 512\;$pixels.
@@ 2.5.1 SourceDetect
The SourceDetect CNN was designed to identify both positive and negative-flux events and to differentiate between them and false objects such as overexposed regions or background artifacts. In the CNN architecture, nine parameters are extracted per detection: the likelihood that it is a real event, its dimensions, its pixel position, and its associated probability for four object classes. These classes are positive detections, negative detections, and two representing false objects such as background artifacts. The model was trained and tested using a dataset of 12,800, $16\times 16$ pixel images with a background generated from a Gaussian distribution to mimic the real reduced TESS images. Each image contained up to two objects from the four classes with varying brightnesses and pixel positions. Further details on SourceDetect will be presented in (Moore et al., in prep.).
@@ 2.5.2 Starfind
As the TESS ePSF is under-sampled, the shape changes significantly under sub-pixel shifts. To account for the variety of appearances for the ePSF we run starfind five times with kernels that correspond to a perfectly centered source, then sources shifted by 0.25 pixels in x and y for each corner. While this combination of sources will not perfectly represent all possible TESS ePSF configurations, it is effective at identifying sources. We compress the results from the 5 searches into a single dataframe, dropping co-located duplicates.
@@
This starfinding detection method outputs a dataframe similar to that produced by SourceDetect. If both SourceDetect and starfind are run on the data, as is default in TESSELLATE, the two dataframes are joined, retaining only the first instance of detections that are spatially located and occur in the same frame.
@@ 2.5.3 Event Isolation
Following the detection of point sources in each TESS FFI, we then group sources into spatially co-located “objects”, which can then have temporally continuous “events” associated with them. For each object, we consider negative and positive detections independently to create negative and positive events. As each event is separated we are able to identify and analyze independent events, such as multiple stellar flares (e.g. Figure 1 flare star example), or identify individual peaks and troughs of variable stars.
@@
At the end of this process we have a data frame containing all independent events and key parameters which can be used for event filtering. For each event we save the pipeline light curve, and generate figures to be used in § 2.9.
@@ 3.1 Detection Efficiency
Understanding the detection efficiency of TESSELLATE is crucial for calculating its expected detection rate of real events. To this end, we inject simulated sources into the datacubes and compute recovery rates for each of TESS’s cameras. These sources are generated by TESS_PRF to match the variable ePSF of TESS, and are randomly assigned an $(x,y)$ position on the detector (at least 3 pixels from the image boundaries), a frame number between 0 and the total number of frames, and a brightness. The brightness is drawn from a log-uniform sample, ranging from 375 to 1.5 TESS counts; using the average TESS zeropoint of 20.44 (Vanderspek et al., 2018), this corresponds to an $m_{T}$ of between $14$ and $20$. We apply both detection methods independently, which allows us to compare their detection efficiencies.
The recovery rates for our test data for Sector 29 are shown in Figure 5. From both methods we have high recovery rates for all cameras until $\sim 17\;m_{T}$ at which point the recovery rate declines significantly. This decline coincides with a source brightness equivalent to a signal to noise ratio $<3$ when considering the average background. As the noise profile is not uniform across cuts, faint sources are only recoverable in areas with low noise.
@@ 3.2 Event Rates
$\left(-16.5\,\rightarrow\,-20\right)$ $\left(-16\,\rightarrow\,-22\right)$ On-axis Off-axis
S29 $0.03^{+0.06}_{-0.02}$ $0.23\pm 0.09$rightarrow $29\pm 11$ $0.004\pm 0.004$rightarrow $15\pm 15$ $5\pm 3$ $14\pm 4$
S01-S78 $2^{+4}_{-2}$ $17\pm 6$rightarrow $2102\pm 788$ $0.3\pm 0.3$rightarrow $1093\pm 1092$ $381\pm 20$ $1026\pm 32$
The recovery rates are shown in Table 2, where we estimate the number of events for Sector 29, and extrapolate to all observed sectors, using the Sector 29 recovery rates. In this simple estimation, our Kilonvae rates are consistent with Mo et al. (2023), however, it is likely that we overestimate the number of events that TESSELLATE will recover. Regardless, this does indicate that TESS is likely to have observed a significant number of known faster transients which currently lie undiscovered in archival TESS data, even if just a single frame is present. As we do save single frame detections, studies on these cases will be conducted once a large population is obtained. Furthermore, after searching each sector of TESS data, we will be able to conduct detailed analysis of event recovery to create strong constraints on the rates of fast extragalactic transients in the local universe.
@@ 4.1 Fast Transients
The primary focus of the TESSELLATE survey is to discover new transients. This designation covers all events which are one-off or irregular in nature, including GRB afterglows, dwarf novae, gravitational microlensing events, kilonovae, and other not-yet-cataloged fast phenomena. Longer-duration transients, such as supernovae, are outside the scope of this survey; while detectable by TESSELLATE, these events are routinely discovered from the ground and exist outside of the unique discovery space for TESS. Therefore, we do not discuss them here.
@@ 4.1.2 Cataclysmic Variables
A significant discovery space still exists for CVs, particularly for high cadence photometry. Even the sample of nearby CVs that are within 150 pc is predicted to be only $77\pm 10$% complete (Pala et al., 2020). Due to the large survey volume and cadence, TESS is an ideal instrument for discovering new out-bursting CVs, such as the previously unidentified CV in Figure 1. With TESSELLATE we will add to the out-bursting CV population with high temporal resolution.
@@ 4.3 Minor Planets
Although TESSELLATE only provides discrete detections of minor planets, we can perform forced photometry to extract lightcurves of all known objects. We stitch together lightcurves of asteroids that cross multiple cuts. The resulting high-cadence lightcurves of minor planets can be used to analyze object morphology and rotation/tumble periods. An example lightcurve segment for the asteroid (5254) Ulysses from Sector 29 is shown in Figure 8. With the Sector 29 light curve we identify a rotation period of $28.71\pm 0.01\;$hours, which is consistent with the asteroid Lightcurve Database (LCDB 2023 release, Warner et al., 2009). Population and individual target analysis with TESSELLATE will be released in upcoming work.
@@ 4.5 Transiting Exoplanets
Figure 9: The light curves of known exoplanet host star WASP-62 (Top Panels) and a previously undiscovered exoplanet candidate (Bottom Panels). The left panels are replotted in their respective right panels after being phase-folded and fit with a standard Gaussian model (red line). Both of these objects were detected by SourceDetect.
Through our untargeted search, we are also able to detect transiting exoplanets. The post-processing pipeline that we discuss in § 2.8 has recovered known transiting exoplanets and produced tens of new candidates just from Sector 29 data. While other pipelines have conducted exoplanet searches of FFI data, they often impose magnitude limits. For example, the search described in Barclay et al. (2018) only considers stars with $I$ magnitudes $\leq 11$ for sun-like stars and $\leq 15$ for M-dwarfs. With TESSELLATE we impost no magnitude limit, and therefore are able to explore a new parameter space for identifying exoplanet candidates of faint stars.
For Cameras 3 and 4 in Sector 29, we recover 8 out of 27 known exoplanet host stars. Due to our search method, all known exoplanet candidates that we recover must have significant dips, therefore, we are unlikely to recover candidates with high contrast between stellar and planetary radii. In Figure 9 we show example light curves of the recovered exoplanet WASP-62 and a TESSELLATE exoplanet candidate.
As many of the host stars for TESSELLATE exoplanet candidates are faint, follow-up and confirmation will be challenging. As with transient detections, we will make all exoplanet candidates public as part of the TESS community target of interest program.
@@ 5 The TESSELLATE Sky Survey
TESS has observed $>90$% of the sky during its six years of operation, with most areas receiving multiple visits. With TESSELLATE, we can build an extensive catalog of variable sources that cover a broad range of astronomy. Following the vetting process outlined in § 2.7, we collate all targets and build the TESSELLATE Sky Survey (TSS).
@@ 6 Conclusion
TESS observations now tile the sky multiple times, enabling an exploration of the rapid time domain. With TESSELLATE, we are able to piece together the variable sky and better understand our dynamic Universe.
H.R. is supported by an Australian Government Research Training Program (RTP) Scholarship. R.R.H., A.M, C.M, and Z.G.L. are supported by the Royal Society of New Zealand, Te Apārangi through the Marsden Fund Fast Start Grant M1255. R.R.H is also supported by the Rutherford Foundation Postdoctoral Fellowship RFT-UOC2203-PD. M.T.B. appreciates support by the Rutherford Discovery Fellowships from New Zealand Government funding, administered by the Royal Society Te Apārangi. We thank the Zooniverse beta testers for Cosmic Cataclysms, in particular the users: Marcossilva, Wirg78, nathancbell, Kutchkutchela, InoSenpai, skraczek, As1310, Mayahn, jklingele, Gammapat, julienlg, remaap24, cesanjerry, captaincat, Barbalbero, andre.andrade, Enriquepa, jrukmini, cosmicTaryn, mprasad, Brbljusa, Tamerciftci; and project translators: InoSenpai, aglr, Michael_Moons-654, Cledison Marcos da Silva. This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government.

### v2

@@ 1 Introduction
The Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2014) has imaged nearly the entire sky in a broad red to near-infrared band since starting operations in July 2018. Its wide field of view (FoV) of 96$\times$24 deg2 and its continuous, high-cadence observation strategy have allowed TESS to fill a niche role in the transient community. While initially designed to hunt for exoplanets in the Milky Way, its science use has since expanded widely, including the study of variable stars (Bruch, 2022; Liu & Qian, 2023; Fetherolf et al., 2023), supernovae (SNe, Fausnaugh et al., 2021; Vallely et al., 2021; Fausnaugh et al., 2023; Wang et al., 2023, 2024), gamma-ray burst (GRB) afterglows (Smith et al., 2021; Fausnaugh et al., 2023; Roxburgh et al., 2023; Jayaraman et al., 2024), asteroids (Pál et al., 2018, 2020; Woods et al., 2021; McNeill et al., 2023), comets (Ridden-Harper et al., 2021a; Farnham et al., 2021), exocomets (Zieba et al., 2019; Dobrycheva et al., 2023), stellar flares (Günther et al., 2020; Pietras et al., 2022), and other transient phenomena.
As the TESS mission has evolved, the volume of data produced from each target field, known as a TESS ‘sector’, has greatly increased. This is due to the progressive shortening of its observation cadence: it began by capturing full-frame images (FFIs) every 30 minutes, shifted to a 10-minute cadence between 5th July 2020 and 1st September 2022, and has since been observing with a cadence of 200 seconds.
The data from each sector is released to the public through the MAST archive (STScI, 2022) as a time-series of FFIs that are ‘calibrated’ but not background subtracted. As such, observers are required to perform their own photometric extraction for sources that are not automatically extracted through the SPOC pipeline (Caldwell et al., 2020). There are a number of successful reduction pipelines available to the public e.g. TESSreduce (Ridden-Harper et al., 2021b), the Quick Look Pipeline (QLP; Fausnaugh et al., 2021; Huang et al., 2020a, b; Fausnaugh et al., 2020; Kunimoto et al., 2021, 2022), eleanor (Feinstein et al., 2019), and DIAmante (Montalto et al., 2020). Each pipeline has their own advantages and disadvantages due to unique choices of techniques. However, these pipelines are often small-scale, focusing on producing cutouts of TESS data around desired coordinates or sources. This means that captured information has been left unexamined in the gaps between targeted searches.
@@
In this paper, we introduce TESSELLATE: the TESS Extensive Lightcurve Logging and Analysis of Transient Events pipeline, the first system capable of automatically searching every TESS frame to document all fast transient events captured by TESS. For this pipeline, we use TESSreduce to remove the scattered light background and create difference images. TESSELLATE is purpose-built for use on the OzSTAR supercomputing node held at Swinburne University, and is wrapped in a user-friendly python package [^1] DOI:https://doi.org/10.5281/zenodo.15954684 that allows for simple but thorough control of the various processing steps.
@@ 2.2 Cutting Datacubes
In this stage, each of the generated datacubes are further divided into a number of cutouts, also completed using the astrocut package. This division enables further parallel processing and decreases the size of individual reduction tasks. Furthermore, the TESSreduce package produces more reliable subtractions of the scattered light background for smaller regions with less dynamic range in the background. The number of cutouts is adjustable; by default each datacube is divided into 16 equal cutouts of $\times 512\;$ 512pixels.
@@ 2.5.1 SourceDetect
The SourceDetect CNN was designed to identify both positive and negative-flux events and to differentiate between them and false objects such as overexposed regions or background artifacts. In the CNN architecture, nine parameters are extracted per detection: the likelihood that it is a real event, its dimensions, its pixel position, and its associated probability for four object classes. These classes are positive events, negative events, and two representing false objects such as background artifacts. The model was trained and tested using a dataset of 12,800, $\times 16$ 16 pixel images with a background generated from a Gaussian distribution to mimic the real reduced TESS images. Each image contained up to two objects from the four classes with varying brightnesses and pixel positions. Further details on SourceDetect will be presented in (Moore et al., in prep.).
@@ 2.5.2 Starfind
As the TESS ePSF is under-sampled, the shape changes significantly under sub-pixel shifts. To account for the variety of appearances for the ePSF we run starfind five times with kernels that correspond to a perfectly centered source, then sources shifted by 0.25 pixels in x and y for each corner. While this combination of sources will not perfectly represent all possible TESS ePSF configurations, it is effective at identifying sources. We compress the results from the 5 searches into a single table, dropping co-located duplicates.
@@
This starfinding detection method outputs a table similar to that produced by SourceDetect. If both SourceDetect and starfind are run on the data, as is default in TESSELLATE, the two tables are joined, retaining only the first instance of detections that are spatially located and occur in the same frame.
@@ 2.5.3 Event Isolation
Following the detection of point sources in each TESS FFI, we then group sources into spatially co-located “objects”, which can then have temporally continuous “events” associated with them. For each object, we consider negative and positive flux detections independently to create negative and positive events. As each event is separated we are able to identify and analyze independent events, such as multiple stellar flares (e.g. Figure 1 flare star example), or identify individual peaks and troughs of variable stars.
@@
At the end of this process we have a table containing all independent events and key parameters which can be used for event filtering. For each event we save the pipeline light curve, and generate figures to be used in § 2.9.
@@ 3.1 Detection Efficiency
Understanding the detection efficiency of TESSELLATE is crucial for calculating its expected detection rate of real events. To this end, we inject simulated sources into the datacubes and compute recovery rates for each of TESS’s cameras. These sources are generated by TESS_PRF to match the variable ePSF of TESS, and are randomly assigned an $(x,y)$ position on the detector (at least 3 pixels from the image boundaries), a frame number between 0 and the total number of frames, and a brightness. The brightness is drawn from a log-uniform sample, ranging from 375 to 1.5 TESS counts; using the average TESS zeropoint of 20.44 (Vanderspek et al., 2018), this corresponds to an $m_{T}$ of between 141414 and 202020. We apply both detection methods independently, which allows us to compare their detection efficiencies.
The recovery rates for our test data for Sector 29 are shown in Figure 5. From both methods we have high recovery rates for all cameras until $\sim 17\;m_{T}$ at which point the recovery rate declines significantly. This decline coincides with a source brightness equivalent to a signal to noise ratio <3<3< 3 when considering the average background. As the noise profile is not uniform across cuts, faint sources are only recoverable in areas with low noise.
@@ 3.2 Event Rates
[…]
S29 $0.03^{+0.06}_{-0.02}$ 0.23±0.090.23$\pm$ 0.090.23 ± 0.09 $\rightarrow$ 29±1129$\pm$ 1129 ± 11 $\pm 0.004$ 0.004 $\rightarrow$ $\pm 15$ 15 5±35$\pm$ 35 ± 3 14±414$\pm$ 414 ± 4
S01-S78 $2^{+4}_{-2}$ 17±617$\pm$ 617 ± 6 $\rightarrow$ 2102±7882102$\pm$ 7882102 ± 788 $\pm 0.3$ 0.3 $\rightarrow$ 1093±10921093$\pm$ 10921093 ± 1092 381±20381$\pm$ 20381 ± 20 1026±321026$\pm$ 321026 ± 32
The recovery rates are shown in Table 2, where we estimate the number of events for Sector 29, and extrapolate to all observed sectors, using the Sector 29 recovery rates. In this simple estimation, our Kilonova rates are consistent with Mo et al. (2023), however, it is likely that we overestimate the number of events that TESSELLATE will recover. Regardless, this does indicate that TESS is likely to have observed a significant number of known faster transients which currently lie undiscovered in archival TESS data, even if just a single frame is present. As we do save single frame detections, studies on these cases will be conducted once a large population is obtained. Furthermore, after searching each sector of TESS data, we will be able to conduct detailed analysis of event recovery to create strong constraints on the rates of fast extragalactic transients in the local universe.
@@ 4.1 Fast Transients
The primary focus of the TESSELLATE survey is to discover new transients. This designation covers all events which are one-off or irregular in nature, including GRB afterglows, classical and dwarf novae, gravitational microlensing events, kilonovae, and other not-yet-cataloged fast phenomena. Longer-duration transients, such as supernovae, are outside the scope of this survey; while detectable by TESSELLATE, these events are routinely discovered from the ground and exist outside of the unique discovery space for TESS. Therefore, we do not discuss them here.
@@ 4.1.2 Cataclysmic Variables
A significant discovery space still exists for CVs, particularly for high cadence photometry. Even the sample of nearby CVs that are within 150 pc is predicted to be only 77±1077$\pm$ 1077 ± 10% complete (Pala et al., 2020). Due to the large survey volume and cadence, TESS is an ideal instrument for discovering new out-bursting CVs, such as the previously unidentified CV in Figure 1. With TESSELLATE we will add to the out-bursting CV population with high temporal resolution.
@@ 4.3 Minor Planets
Although TESSELLATE only provides discrete detections of minor planets, we can perform forced photometry to extract lightcurves of all known objects. We stitch together lightcurves of asteroids that cross multiple cuts. The resulting high-cadence lightcurves of minor planets can be used to analyze object morphology and rotation/tumble periods. An example lightcurve segment for the asteroid (5254) Ulysses from Sector 29 is shown in Figure 8. With the Sector 29 light curve we identify a rotation period of 28.71±0.0128.71$\pm$ 0.01\;28.71 ± 0.01hours, which is consistent with the asteroid Lightcurve Database (LCDB 2023 release, Warner et al., 2009). Population and individual target analysis with TESSELLATE will be released in upcoming work.
@@ 4.5 Transiting Exoplanets
Figure 9: The light curves of known exoplanet host star WASP-62 (Top Panels) and a previously undiscovered exoplanet candidate TSS J0547-6921 (Bottom Panels). The left panels are replotted in their respective right panels after being phase-folded and fit with a standard Gaussian model (red line) which is used in the TESSELLATE exoplanet vetting criteria.
Through our untargeted search, we are also able to detect transiting exoplanets. The post-processing pipeline that we discuss in § 2.8 has recovered known transiting exoplanets and produced tens of new candidates just from Sector 29 data. While other pipelines have conducted exoplanet searches of FFI data, they often impose magnitude limits. For example, the search described in Barclay et al. (2018) only considers stars with $I$ magnitudes 11$\leq$ 11≤ 11 for sun-like stars and 15$\leq$ 15≤ 15 for M-dwarfs. With TESSELLATE we impost no magnitude limit, and therefore are able to explore a new parameter space for identifying exoplanet candidates of faint stars.
For Cameras 3 and 4 in Sector 29, we recover 8 out of 27 known exoplanet host stars. Due to our search method, all known exoplanet candidates that we recover must have significant dips, therefore, we are unlikely to recover candidates with high contrast between stellar and planetary radii. In Figure 9 we show example light curves of the recovered exoplanet WASP-62 and a TESSELLATE exoplanet candidate TSS J0547-6921.
As many of the host stars for TESSELLATE exoplanet candidates are faint, follow-up and confirmation will be challenging. As with transient detections, we will make all exoplanet candidates and eclipsing binaries public as part of the TESS community target of interest program.
@@ 5 The TESSELLATE Sky Survey
TESS has observed >90>90> 90% of the sky during its six years of operation, with most areas receiving multiple visits. With TESSELLATE, we can build an extensive catalog of variable sources that cover a broad range of astronomy. Following the vetting process outlined in § 2.7, we collate all targets and build the TESSELLATE Sky Survey (TSS).
@@ 6 Conclusion
TESS observations now tile the sky multiple times, enabling an exploration of the rapid time domain. This coverage will only improve with TESS’s Extended Mission 3 (EM3), which is expected to begin in September 2025. EM3 will increase the temporal baseline with longer sectors with significant overlap at mid-latitudes. TESSELLATE is perfectly poised to take advantage of this extended monitoring, piecing together the variable sky to better understand our dynamic Universe.
H.R. is supported by an Australian Government Research Training Program (RTP) Scholarship. R.R.H., A.M, C.M, and Z.G.L. are supported by the Royal Society of New Zealand, Te Apārangi through the Marsden Fund Fast Start Grant M1255. R.R.H is also supported by the Rutherford Foundation Postdoctoral Fellowship RFT-UOC2203-PD. M.T.B. appreciates support by the Rutherford Discovery Fellowships from New Zealand Government funding, administered by the Royal Society Te Apārangi. The material is based upon work supported by NASA under award number 80GSFC24M0006. We thank the Zooniverse beta testers for Cosmic Cataclysms, in particular the users: Marcossilva, Wirg78, nathancbell, Kutchkutchela, InoSenpai, skraczek, As1310, Mayahn, jklingele, Gammapat, julienlg, remaap24, cesanjerry, captaincat, Barbalbero, andre.andrade, Enriquepa, jrukmini, cosmicTaryn, mprasad, Brbljusa, Tamerciftci; and project translators: InoSenpai, aglr, Michael_Moons-654, Cledison Marcos da Silva. This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government.